Victor would like to buy his first car and the one he has his eye on is $25,000 plus an extra 13% HST for a total price of $28,250. The dealership has a deal for 0% down payment and charges 2.89% interest on the loan. Victor plans to make car loan payments weekly and has accepted the maximum loan repayment period of 8 years.
How much will his weekly care loan payment be?
How much will he have paid to the dealership by the time his loan is paid off?
How much interest will be paid?

Answer by addingup(3677)

So Victor is financing $28,250, correct?
8*52 = 416 payments
at an interest rate of 2.89% per annum (you didn't say, interest rates are usually per year? That's the assumption I made here, if it's not you'll have to either change the numbers)
-- - - - - - - - - - - - - - - - - -
A = P*(r(1+r)^n)/(((1+r)^n)-1)
A = 28250(0.00055577(1+0.00055577)^416)/(((1+0.00055577)^416)-1)
A = 19.783115/0.26
A = 76.089 you can round this to 76.10. This is the amount of the weekly payment.
.
76.089*416 = $31,653 this is the amount he will have paid the dealership
31,653-28,250 = 3,403 this is the amount of interest